# GRAND: Graph Release with Assured Node Differential Privacy

### v2

@@ 6.1 Results under the inner product latent space model
Table 1: Node-level statistic distribution preservation results with $n=2000$ under the inner product latent space model, measured by the average Wasserstein distance over 100 replications. The numbers in the parentheses are the corresponding standard errors.
Metric $d$ $\varepsilon$ $\rho=0.025$ $\rho=0.05$ $\rho=0.1$
Hat GRAND Laplace Hat GRAND Laplace Hat GRAND Laplace
Node Degree 3 1 0.047 (<0.001<0.001) 0.068 (0.003) 3.033 (0.002) 0.017 (<0.001<0.001) 0.05 (0.002) 2.334 (0.001) 0.008 (<0.001<0.001) 0.031 (0.002) 1.632 (0.001)
2 0.047 (<0.001<0.001) 0.063 (0.002) 2.95 (0.002) 0.017 (<0.001<0.001) 0.047 (0.002) 2.264 (0.002) 0.008 (<0.001<0.001) 0.031 (0.002) 1.582 (0.001)
5 0.047 (<0.001<0.001) 0.046 (0.001) 2.554 (0.004) 0.017 (<0.001<0.001) 0.036 (0.002) 1.924 (0.004) 0.008 (<0.001<0.001) 0.026 (0.001) 1.333 (0.003)
10 0.047 (<0.001<0.001) 0.047 (0.001) 1.804 (0.006) 0.017 (<0.001<0.001) 0.03 (0.001) 1.3 (0.005) 0.008 (<0.001<0.001) 0.024 (0.001) 0.881 (0.004)
6 1 0.051 (<0.001<0.001) 0.074 (0.002) 3.043 (0.001) 0.019 (<0.001<0.001) 0.051 (0.002) 2.347 (0.001) 0.008 (<0.001<0.001) 0.029 (0.002) 1.642 (0.001)
2 0.051 (<0.001<0.001) 0.068 (0.002) 3.022 (0.001) 0.019 (<0.001<0.001) 0.046 (0.002) 2.332 (0.001) 0.008 (<0.001<0.001) 0.03 (0.001) 1.63 (0.001)
5 0.051 (<0.001<0.001) 0.053 (0.001) 2.895 (0.002) 0.019 (<0.001<0.001) 0.038 (0.002) 2.227 (0.002) 0.008 (<0.001<0.001) 0.029 (0.001) 1.555 (0.001)
10 0.051 (<0.001<0.001) 0.055 (0.002) 2.534 (0.004) 0.019 (<0.001<0.001) 0.033 (0.001) 1.928 (0.003) 0.008 (<0.001<0.001) 0.021 (0.001) 1.336 (0.002)
V-shape 3 1 0.099 (0.001) 0.142 (0.005) 6.134 (0.003) 0.034 (0.001) 0.102 (0.005) 4.7 (0.003) 0.016 (<0.001<0.001) 0.063 (0.003) 3.277 (0.002)
2 0.099 (0.001) 0.132 (0.005) 5.968 (0.004) 0.034 (0.001) 0.097 (0.005) 4.559 (0.004) 0.016 (<0.001<0.001) 0.063 (0.003) 3.178 (0.003)
5 0.099 (0.001) 0.097 (0.003) 5.171 (0.008) 0.034 (0.001) 0.074 (0.003) 3.877 (0.008) 0.016 (<0.001<0.001) 0.053 (0.003) 2.677 (0.006)
10 0.099 (0.001) 0.097 (0.003) 3.662 (0.012) 0.034 (0.001) 0.061 (0.003) 2.624 (0.01) 0.016 (<0.001<0.001) 0.049 (0.002) 1.772 (0.008)
6 1 0.108 (0.001) 0.155 (0.004) 6.151 (0.003) 0.04 (0.001) 0.103 (0.004) 4.724 (0.002) 0.017 (<0.001<0.001) 0.058 (0.003) 3.298 (0.002)
2 0.108 (0.001) 0.142 (0.004) 6.11 (0.003) 0.04 (0.001) 0.094 (0.003) 4.693 (0.002) 0.017 (<0.001<0.001) 0.06 (0.003) 3.274 (0.002)
5 0.108 (0.001) 0.112 (0.002) 5.855 (0.004) 0.04 (0.001) 0.078 (0.003) 4.484 (0.003) 0.017 (<0.001<0.001) 0.058 (0.003) 3.123 (0.003)
10 0.108 (0.001) 0.115 (0.003) 5.131 (0.008) 0.04 (0.001) 0.066 (0.002) 3.885 (0.006) 0.017 (<0.001<0.001) 0.043 (0.002) 2.683 (0.005)
Triangle 3 1 0.055 (0.001) 0.127 (0.006) 8.774 (0.005) 0.029 (0.001) 0.132 (0.007) 6.768 (0.004) 0.015 (<0.001<0.001) 0.094 (0.005) 4.799 (0.003)
2 0.055 (0.001) 0.12 (0.005) 8.608 (0.006) 0.029 (0.001) 0.126 (0.008) 6.622 (0.005) 0.015 (<0.001<0.001) 0.092 (0.005) 4.694 (0.004)
5 0.055 (0.001) 0.122 (0.006) 7.789 (0.009) 0.029 (0.001) 0.1 (0.006) 5.899 (0.009) 0.015 (<0.001<0.001) 0.08 (0.005) 4.142 (0.007)
10 0.055 (0.001) 0.142 (0.006) 6.137 (0.015) 0.029 (0.001) 0.09 (0.005) 4.449 (0.013) 0.015 (<0.001<0.001) 0.077 (0.004) 3.025 (0.01)
6 1 0.066 (0.001) 0.129 (0.005) 8.794 (0.005) 0.034 (0.001) 0.102 (0.006) 6.781 (0.003) 0.017 (<0.001<0.001) 0.073 (0.005) 4.78 (0.003)
2 0.066 (0.001) 0.139 (0.006) 8.754 (0.005) 0.034 (0.001) 0.092 (0.005) 6.75 (0.003) 0.017 (<0.001<0.001) 0.077 (0.005) 4.754 (0.003)
5 0.066 (0.001) 0.163 (0.007) 8.501 (0.005) 0.034 (0.001) 0.085 (0.005) 6.531 (0.005) 0.017 (<0.001<0.001) 0.075 (0.004) 4.592 (0.004)
10 0.066 (0.001) 0.224 (0.009) 7.754 (0.009) 0.034 (0.001) 0.084 (0.004) 5.89 (0.007) 0.017 (<0.001<0.001) 0.057 (0.003) 4.101 (0.006)
Eigen Centrality 3 1 0.012 (0.001) 0.042 (0.003) 0.333 (0.005) 0.01 (0.001) 0.041 (0.003) 0.25 (0.005) 0.011 (0.001) 0.031 (0.002) 0.164 (0.004)
2 0.012 (0.001) 0.039 (0.003) 0.288 (0.005) 0.01 (0.001) 0.04 (0.003) 0.199 (0.005) 0.011 (0.001) 0.028 (0.002) 0.106 (0.003)
5 0.012 (0.001) 0.042 (0.003) 0.198 (0.005) 0.01 (0.001) 0.041 (0.003) 0.108 (0.004) 0.011 (0.001) 0.031 (0.002) 0.043 (0.001)
10 0.012 (0.001) 0.039 (0.004) 0.069 (0.004) 0.01 (0.001) 0.039 (0.003) 0.048 (0.002) 0.011 (0.001) 0.028 (0.002) 0.097 (0.003)
6 1 0.016 (0.001) 0.045 (0.003) 0.46 (0.007) 0.012 (0.001) 0.04 (0.003) 0.408 (0.007) 0.012 (0.001) 0.032 (0.002) 0.332 (0.006)
2 0.016 (0.001) 0.048 (0.003) 0.361 (0.006) 0.012 (0.001) 0.04 (0.003) 0.293 (0.006) 0.012 (0.001) 0.033 (0.002) 0.226 (0.005)
5 0.016 (0.001) 0.048 (0.003) 0.242 (0.005) 0.012 (0.001) 0.042 (0.003) 0.168 (0.005) 0.012 (0.001) 0.031 (0.002) 0.101 (0.004)
10 0.016 (0.001) 0.053 (0.004) 0.145 (0.005) 0.012 (0.001) 0.036 (0.003) 0.066 (0.003) 0.012 (0.001) 0.032 (0.002) 0.045 (0.002)
Harmonic Centrality 3 1 13.597 (0.107) 14.552 (0.604) 579.684 (0.429) 3.616 (0.047) 6.631 (0.292) 453.333 (0.282) 0.691 (0.011) 2.933 (0.151) 392.163 (0.241)
2 13.597 (0.107) 13.014 (0.594) 551.887 (0.617) 3.616 (0.047) 6.335 (0.304) 427.719 (0.508) 0.691 (0.011) 2.927 (0.153) 372.992 (0.463)
5 13.597 (0.107) 7.125 (0.394) 433.659 (0.939) 3.616 (0.047) 4.681 (0.21) 318.755 (0.98) 0.691 (0.011) 2.466 (0.144) 284.812 (0.781)
10 13.597 (0.107) 6.481 (0.264) 276.721 (1.028) 3.616 (0.047) 3.376 (0.17) 171.327 (0.916) 0.691 (0.011) 2.299 (0.116) 157.138 (0.854)
6 1 15.681 (0.1) 15.535 (0.573) 581.975 (0.328) 3.736 (0.038) 6.732 (0.253) 458.477 (0.148) 0.716 (0.01) 2.575 (0.148) 398.107 (0.118)
2 15.681 (0.1) 13.531 (0.593) 574.048 (0.364) 3.736 (0.038) 6.134 (0.212) 452.14 (0.198) 0.716 (0.01) 2.683 (0.14) 392.887 (0.169)
5 15.681 (0.1) 8.566 (0.413) 528.079 (0.618) 3.736 (0.038) 4.812 (0.207) 411.398 (0.502) 0.716 (0.01) 2.589 (0.118) 361.959 (0.389)
10 15.681 (0.1) 7.184 (0.321) 419.364 (0.942) 3.736 (0.038) 3.472 (0.123) 313.355 (0.816) 0.716 (0.01) 1.94 (0.084) 282.385 (0.741)
@@
  • When focusing on the proposed GRAND method, it can be seen that as $\varepsilon$ increases, the performance becomes better for all statistics, which is expected due to the potential privacy-utility tradeoff.
  • Comparing our method with the Hat method, we can see that it indeed yields worse preservation. This is also expected since this reflects the cost of achieving privacy. Noticeably, for weak privacy protection, such as $\varepsilon=10$, the performance of our method becomes comparable to the non-private Hat method. Recall that the Hat method is based on an estimation of a size-$m$ network. This verifies our previous statement that the node-wise estimation basically gives the accuracy as a standard estimation of size-$m$ network. It is also noted that for several metrics, our method gives even a slightly better preservation than the Hat method for $\varepsilon=10$. This is indeed because in our method, we also use the information that the latent vectors are i.i.d following $F$, in processing the network. However, for the Hat method using the estimation of Ma et al. (2020), this distributional information is not used. In other words, when $\varepsilon$ becomes large, the advantage of using the distributional information outweighs the small loss due to the privacy protection.
  • Comparing our method with the Laplace mechanism, where both methods satisfy the same strictness of node DP, it is evident that our method achieves much better preservation of network properties. The Laplace mechanism blindly introduces the noises but fails to retain the network structures.
  • The aforementioned patterns are consistent across different configurations of latent dimensions, network sizes and densities. This demonstrates the effectiveness and robustness of our method across a wide range of scenarios.

### v3

@@ 6 Simulation Experiments
Figure 2: Comparison of node-level statistic distribution preservation for $\rho=0.05$ with $n=4000$ and $d=3$ under the inner product latent space model (LSM) and random dot product graph (RDPG) model. Values are the average Wasserstein distance over 100 replications. The average distances are more than ten times larger than their corresponding standard errors; therefore, standard errors are omitted for brevity.
Metric $\varepsilon$ LSM RDPG
Hat GRAND Laplace Hat GRAND Laplace
Node Degree 1 0.010 0.030 2.340 0.009 0.019 2.053
2 0.010 0.030 2.276 0.009 0.021 1.430
5 0.010 0.023 1.949 0.009 0.018 0.535
10 0.010 0.018 1.327 0.009 0.016 0.487
V-shape Count 1 0.020 0.061 4.696 0.018 0.038 4.121
2 0.020 0.060 4.568 0.018 0.042 2.872
5 0.020 0.047 3.913 0.018 0.037 1.080
10 0.020 0.036 2.666 0.018 0.031 1.005
Triangle Count 1 0.013 0.081 6.768 0.019 0.047 6.532
2 0.013 0.079 6.633 0.019 0.050 5.117
5 0.013 0.066 5.933 0.019 0.042 2.091
10 0.013 0.054 4.491 0.019 0.035 1.385
Eigen Centrality 1 0.007 0.028 0.257 0.020 0.023 0.125
2 0.007 0.029 2.080 0.020 0.023 0.178
5 0.007 0.028 0.115 0.020 0.022 0.299
10 0.007 0.029 0.041 0.020 0.023 0.305
Harmonic Centrality 1 2.046 4.445 884.790 1.162 2.385 659.590
2 2.046 4.301 836.624 1.162 2.558 345.161
5 2.046 3.298 621.660 1.162 2.300 133.620
10 2.046 2.374 324.391 1.162 1.975 124.544
@@
  • When focusing on the proposed GRAND method, it can be seen that as $\varepsilon$ increases, the overall performance becomes better, following the potential privacy-utility tradeoff.